Investors require a return of 14.8% per year to hold a stock. The stock currently does not pay any dividends and is expected to begin paying a dividend of $1.65 in 3 years from now, then the dividends are expected to grow forever at a constant rate of 8.0%. What’s the stock price?

A)

$15.72

B)

$16.05

C)

$18.41

D)

$17.34

E)

$13.74
Answer #1

Value in 2 years = D3/ (Required rate - growth rate)

Value in 2 years = 1.65/ (14.8% - 8%)

Value in 2 years = $24.26

Value today = 24.26/ (1.148)2

Value today = $18.41